OpenStudy (anonymous):

Brooklyn has a goal to save $8,000 to buy a new entertainment system. In order to meet that goal, she deposited $4,132.79 into a savings account. If the account has an interest rate of 4.8% compounded quarterly, approximately when will Brooklyn be able to make the purchase?

OpenStudy (anonymous):

\[FC=IC(\frac{i}{100}+1)^{n}\] Where FC is the goal meet $8000 IC is the inicial capital $4132.79 i the interest every 3 months 4.8% and n is the number of quarters (groups of 3 months ) that are needed. So we must take logaritms in both sides of te equation\[\ln FC =\ln( IC+{\frac{i}{100})^{n}\] Using the rules of logaritms we can obtain \[n=\frac {\ln FC-\ln IC}{\ln(\frac{i}{100}+1)}\] So \[n=\frac {\ln 8000 - \ln 4132.79}{\ln (\frac {4.8}{100}+1)}\] n=14.04 quartes n=14.04 /4=3.51 years n=14.04 x 3= 42.12 months n=42.12 x 30 = 12136.6 días

OpenStudy (anonymous):

The unseen equation is this So we must take logaritms in both sides of te equation \[\ln FC =\ \ln {( IC+\ \frac {i} {100} )^{n}}\]